IMACS progress report no. 54
B. Bigelow 3 October 2002
I) Project Status
Currently scheduled tasks: Status:
a) DSS drive design (Hare) Done
b) DSS drafting (Cleary) Done
c) DSS assembly/testing (Hare, Kowal, Castillo) IP
d) DSS Grating Tilt Mech testing (Hare) IP
e) Calibration system mech. des. (Hare) on hold
f) CCD array assembly (Thompson/Burley) IP
g) Grism procurement (Sutin) IP
h) Grism cell design (Bigelow) Done
i) Grism cell drafting (Cleary) IP
j) Entrance hatch design (Bigelow) Done
k) Entrance hatch drafting (Cleary) Done
l) Entrance hatch fabrication (Kowal) ready to start
m) Short Camera shutter fabrication (Kowal) Done
n) Short Camera shutter assembly (Hare) ready to start
o) Instrument electronics assembly (Bagish) on sched.
p) Control software design (Birk) on sched.
q) Short Camera opto-assembly (Bigelow) IP
r) Instrument enclosures fab (Gunnels/Paragon) Done
s) Instrument enclosures installation (Gunnels) IP
t) Instrument utility wrap fabrication (Rettig) IP
u) IFU design and build (U. Durham) IP
Project events in September:
- First CCD array assembly with all 8 chips
- First spectra with grating, detectors, long camera, and software
- Short camera aspheres approaching completion
Latest project images:
Final Assy: http://www.ociw.edu/instrumentation/imacs/devel/final/index.html
Dewar Assy: http://www.ociw.edu/instrumentation/imacs/devel/dewar/index.html
II) Optics (Sutin)
* The filters are all mounted in their respective labeled frames by
their corners except for the long camera spectrographic dummy filter,
which is still inside the long camera. They need to be RTV'ed in.
* Tinsley has finished S08, and probably finished S01. We await
final S01 data. We hope to sign off later this week.
* The S04 replacement is still at TORC, just going into polish.
Should be finished in 3-4 weeks.
* 300 l/mm grism is at ThermoRGL, having the ruling applied. While
there, the coating was damaged, probably from no fault of their
own, but because the coating failed the "scotch tape test", or
came off of the surface too easily. The coating needs to be
* 600 l/mm grism glass is coated and at HJOL, having the two parts
cemented together. The original choice of cement was too viscous
and dried to fast for a part this large, so we have switched to
a different cement.
* The echellette grating has been requisitioned. The prism detail
design has not been started.
* The 200-micron pinhole for making S-H calibration frames is here.
III) Electronics (Bagish)
- Chassis 2 upgraded with Altera hardstop logic. All chassis upgrades are complete.
- Set up 20 RTD thermal elements for test. Tested o.k.
- All MCS panels have been surfaced and anodized. Four panels are at the engravers.
- Eight cables assembled. 70% completion.
- Installed Disperser Server Tilt A micro-stepping driver.
- Fluid/Air pressure/temperature sensors came in.
- Attach instrument end manifold and wire up individual subsections.
- Engrave remaining panels.
- When installed, measure length of cable wrap and complete assembly of remaining
ends of wrap cables.
- Cable assembly to 90% completion.
IV) Software (Birk/Clardy)
- Gratings/Disperser Wheel: No progress (under construction) For
October I plan to continue testing of the grating mechanism
(ie. control-loop) and start testing the disperser wheel.
- SH-, PR- and CF-Guiders: Conducted flexure tests with AlanD and
Tyson. I tested and fine-tuned the control loops for the theta axis
of the PR- and SH-guiders. Improved resolution for the theta
drive(s): The ratio of the current system is 1.3 encoder-steps per
motor-step. Ie. increasing the encoder resolution (from 5um to 1
um) requires a change of the motor-gear(s).
- DSP/Science Array Software First data acquisition using the IMACS
software suite (CCDServer, CamGUI and QlTool) reading one chip at a
- HP-Sensor Unit Final test of 20 temperature sensors was
successful. All sensors are within +-0.06 deg. Celsius.
- 2nd shutter no progress (not assembled)
- Mask Server For October I plan to test the new version with 8 masks
- Got the changes to GMAP made by Skip merged with those made to
support IMACS guider fields. Verified that features of all these
changes are working.
- Looked into changing how the stars are selected for the high
density USNO-A catalog. There is a great difference in different
parts of the sky in star density, so one magnitude limit based on
the average density does not always work, and can suppress the
plotting of suitable guide stars in sparse areas.
- I made an interactive program to test the optics transform
polynomials supplied by Brian. The optics code will be needed in
the GMAP program because the guide stars are significantly
displaced from a linear approximation when we have a guider field
15 arc minutes off axis.
- I am starting to outline the operation and appearance of the
Astronomer's GUI, which will be a purpose-built editor for the mask
generation program's observation file. It will present choices and
defaults, and check input for consistency.
V) Detector Systems (Thomson/Burley) (Bigelow reporting)
1) Getting the fiber link to work (prereq for the 8-chip readout)
2) Get all detectors working together in dewar
3) Burley in Pasadena: 10/1 - 10/12
VI) Mechanics (Bigelow/Hare)
1) Continued short camera assembly
2) Finished instrument hatch/calibration screen design, drafting done
3) Finished grism cell designs, drafting started
1) Continue short camera opto-assembly
2) Alignment and testing of slit mask stability
3) Grism cell fab drawings to shop
4) Order final cryotiger hose assemblies for both cameras
5) Finish plumbing design for guider cameras
6) Return flexure control stage to PI for extended cables and re-calibration
7) Progress revised grating cell and tilt mechanism designs
* Continued opto-mechanical and electro-mechanical assembly and testing of all Guide
* Designing counterbalance for the Shack-Hartmann Guide Camera.
* Assembly and testing of the Grating Tilt Mechanism. Investigating flexure issues.
* Ordered parts for the DSS Drive Assembly and Cable Wrap.
* Fabrication DSS Drive Assembly.
* Continue assembly, testing, and rework of all guide cameras as necessary.
* Continued testing of the Grating Tilt Mechanism.
* Assembly and testing of the DSS Drive Assembly.
* Assembly and testing of the Cable Wrap Subassembly.
* Assembly of the second shutter.
* Resume design work of the Calibration Unit.
VII) Fabrication (Kowal)
1) Shutter v2.0 is back from being teflon anodized.
2) Machined 6 new pillars for the G.T.M.
3) Four stainless steel G.T.M. base plates have been finished.
4) Made new titanium brace for G.T.M. Trunion.
5) Started machining Carousel Drive Base.
Expectations for October:
1) Finish Carousel Drive Assembly
2) Finish Hatch
3) Finish remainder of new Imaging Mirror Cell.
4) Start Machining Grism Cells.
- AOA microlens arrays arrived and tested thoroughly:
Input arrays: - Back focal length: Mean value 18.5 microns
- Pitch: 402.375 microns (goal 402.5 microns - acceptable)
- 15.75 out of 1320 lenses wrong, mostly at edge where not used.
- Very smooth and clean surfaces, well defined diff-limited foci
Output arrays: - Back focal length: 178 +/- 15 microns mean value (acceptable)
- Pitch: 244.957 +/- 0.039 microns (acceptable)
- Faulty lenslet statistics: 5 % of lenses show strong damages.
There are scratchy, rough surfaces on many surfaces:
Coverage 1/2 to 1 lens: 14 %
Coverage 1/4 to 1/2 lens: 28 %
Coverage 1/8 to 1/4 lens: 17 %
Coverage < 1/8 lens: 43 %
- The arrays are badly positioned on the substrates, some
inclined and all in a position wrong by 5 mm.
- About half the batch is quite dusty.
The decision is to reject the whole output array batch. This will
result in a delay estimated to be 6 weeks for re-manufacturing. Even
when the errors were accepted the misreading of our drawing results in
difficulties to get the lenslet arrays mounted.
- Fiber bundles arrived at Durham. Pitch was measured to 244.1 microns
(goal 244.3), which is acceptable.
- Base plate two (final version), comb and auxiliary parts anodized.
- V-rails hard anodized.
- Fiber glass housing: Delayed. Company damaged perform and has to rework it. It is
planned to visit the company soon to assure that they can overcome this problems.
- Completion of volume dummy is still expected after delivery of the fiber glass
Procurement: Micro lens coating delayed by output array acceptance problem.
Integration and Test:
- Alignment document preparation is going on.
- Micro lens test setup works sufficiently. Additionally, some PSF
measurements of the poor output arrays are made with high dynamic
range to monitor the consequences on performance. Furthermore a
white light interferometer will be used to quantify the surface
- Large alignment and assembly jig is finished. Hence, all but some
minor parts for alignment and installation are ready.
IX) Other Business
Next meeting: 10 AM, 11/6/02